# AllSpark[Uncaptioned image]: A Multimodal Spatio-Temporal General Intelligence Model with Thirteen Modalities

### v2

@@ III-A Theory
Inspired by our observation, to balance cohesion and autonomy among multiple modalities, we introduce the fundamental principle of Language as Reference Framework (LaRF).
Figure 2: Guided by the LaRF principle, multimodal data are transformed into a token-context structure similar to language under their respective prior assumptions while maintaining the autonomy of the multimodalities and achieving cohesion among modalities, thus enabling interpretation of multimodal data in the unified language representation space.
In terms of the cohesion of multimodalities, the high heterogeneity between multiple spatio-temporal modalities is a major challenge, while the LaRF principle defines the alignment anchor between multimodalities as language explicitly. As shown in Figure 2, we observe that language is encoded by tokens and their contexts, and this structure can be extended to most spatio-temporal modalities. Therefore, we can align highly heterogeneous spatio-temporal modalities to language modalities in structure and semantics, enabling multimodal interpretation in the unified representation space of language. In addition, the important guiding role of natural language prompts is also one of the reasons why the LaRF principle can achieve cohesion among modalities.
@@
    Interpretability: Deep learning methods have long been characterized as opaque ”black boxes.” However, a multimodal intelligent system constructed based on the LaRF can directly leverage language as a tool. This facilitates the output of interpretable reasoning chains that humans can understand, thereby achieving true explainable artificial intelligence.
@@
Figure 3: AllSpark Architecture. Multimodal data are extracted by their respective modal encoders into token sequences. After undergoing dimension alignment with modal-specific text prompt tokens through a modal bridge, the text prompt tokens and modal tokens enter a large language multimodal model for interpretation. Finally, the interpretation result is matched with downstream tasks by task-specific task heads.
@@ III-B Overview
TABLE I: List of symbols
@@
Symbol
Meaning
Symbol
Meaning
$M$
AllSpark
$m_{i}$
modal i
$f_{i}$
modal encoder for $m_{i}$
$q$ query vectors of the bridge
$\Phi$
modal bridge
$s_{i}$ tokens of $m_{i}$
$T$
text tokenizer
$p_{i}$ text prompt for $m_{i}$
$F$
Multimodal LLM
$\theta$ model parameters
$H_{task}$
Task Head
$y$ label
$L$
loss function
$t_{i}$ tokens of $m_{i}$
${Enc}_{n}$
n-layer transformer encoders
$W$ weights of the linear layer
$Emb$
embedding layer
$w_{i}$ word of the code or text
${Enc}_{ResNet}$
ResNet
$Sparse$
sparse module of the TubeViT
$Grouper$
point grouper of the PointBERT
$Conv1d$
1D convolution layer
$FFN$
feedforward network
$\sigma$
softmax
Guided by the principles mentioned above, the AllSpark we proposed consists of five modules: the modal encoder, modal bridge, text tokenizer, multimodal large language model, and task head. The overall architecture is depicted in Figure 3.
@@
Here, $\{m_{RGB},m_{MSI},m_{HSI},\ldots,m_{i},\ldots,m_{Video}\}$ represents inputs from various modalities, where $m_{RGB}$ represents the RGB modality, $m_{MSI}$ represents the MSI modality, and so on. $f_{i}$ is defined as the modal encoder for the $m_{i}$ modality, $\Phi$ represents the modal bridge, and $q\in R^{N*D}$represents $N$ learnable vectors of dimension $D$ in the modality bridge, where $D$ is set to 4096, representing the dimensionality of the multimodal large language model. The input data $m_{i}$ of each modality are mapped to a token sequence $s_{i}\in R^{N*D}$ of the same dimension as the language model. Table I summarizes the main mathematical symbols and their meanings in the model.
To achieve cohesion among modalities, we employ a unified multimodal large language model to parse data from various modalities. The text tokenizer and multimodal large language model in AllSpark are based on the visual-language model Lynx[10]. To extend Lynx to thirteen spatio-temporal modalities, we designed specific text prompts for each modality to guide the model in correctly parsing information from each modality. Additionally, Lynx incorporates several lightweight multimodal adapter layers internally to accommodate multimodal inputs. We continue this design and do not freeze the parameters of the adapter layers during training to enhance the model’s adaptability to other spatio-temporal modalities. Finally, we acknowledge that a gap exists between the parsing results of the model and those of the downstream task. Therefore, we design specific task heads for each task to enhance the model’s generalization capability.
@@ III-C Independent Encoder for Each Modality
Code & Text: Code is essentially a specialized form of language, but due to its distinct properties such as having a strict syntax, being unambiguous, and being capable of interacting with machines. we separate it as a distinct modality. Exploring intelligent methods for the code modality is crucial for reliable AI reasoning and achieving interaction between intelligent models and the real world. Since the Lynx model is a language model, we do not design an additional modal encoder for the code and text modalities. Instead, we directly utilize Lynx’s text tokenizer, i.e., $f\left(m_{Text/Code}\right)=T\left(m_{Text/Code}\right)$, where $m_{Text/Code}=\{w_{1},w_{2},w_{3},\ldots\}$ represents the sequence of words in the text or code.
@@
RGB: RGB imagery represents the visible light spectrum and reflects the electromagnetic characteristics of objects that emit or reflect visible light waves. It is the most common modality used in visual research. RGB imagery is a standard three-band image, i.e., $m_{RGB}\in R^{H*W*3}$. For the modal encoder of this modality, we adopted the visual encoder from the Lynx model: the EVA. The EVA is a large visual backbone model composed of 40 stacked transformer blocks with a width of 1408. During the experiments, AllSpark loaded the official weights of the EVA model and froze them during training.
MSI: Multispectral imagery is a modality extensively studied in the remote sensing field. Typically, it includes more than the three visible light bands present in the RGB modality. MSIs incorporate multiple nonvisible light bands, such as near-infrared, shortwave infrared, coastal atmospheric aerosol, and cirrus bands. Therefore, the number of channels in the imagery is usually greater than the three bands in RGB imagery, i.e., $m_{MSI}\in R^{H*W*C}$, where $C>3$. We extended the PatchEmbed of the standard ViT[18], modifying its channel count to match the number of bands in the input multispectral imagery. This adaptation allows it to serve as the feature encoder for the multispectral modality.
@@
Table: The table modality consists of several rows and columns, where each row represents a record and each column represents a different attribute, i.e., $m_{Table}\in R^{{row}*{col}}$. The modal encoder for the table modality inherits the design from TabFormer[36]: first, based on the different degrees of discreteness for each column attribute, we use independent embedding layers to encode discrete and continuous values separately. Subsequently, we employ a single-layer transformer encoder to further extract features. The entire process can be formalized as $f\left(m_{Table}\right)={Enc}_{1}\left(Emb(m_{Table})\right)$.
@@
Infrared: The dataset we selected for infrared detection was SYSU-MM01, which is, in fact, a multimodal reidentification dataset comprising both visible and near-infrared (NIR) modalities. The model needs to simultaneously process infrared and visible light images, i.e., $m_{Infrared}=\{{I_{r}\in R}^{H*W*3},{I_{i}\in R}^{H*W*3}\}$. We inherited and slightly modified the design of DDAG[38], discarding the parameters of the shared ResNet-50 layers in DDAG and employing two independently parameterized ResNet-50 networks. These networks extract features from infrared and visible light images, and this process can be formalized as $f\left(m_{Infrared}\right)={Enc}_{ResNet}\left(I_{r}\right)\oplus\ {Enc}_{ResNet}\left(I_{i}\right)$.
Graph: A graph is constructed from a series of nodes and edges, where the attributes of the nodes and the adjacency attributes of the nodes reflect the majority of the features of the graph, i.e., $m_{Graph}\in R^{K*d}$, where K is the number of nodes and d is the node feature dimension. In AllSpark, the modality encoder of a graph is based on the STAEformer[39], whose main design idea is to first use a linear layer to extend its feature dimension and then use several embedding layers to separately encode features such as the node’s characteristics, spatial characteristics, and temporal characteristics. The entire process can be formalized as $f\left(m_{Graph}\right)={Emb}_{node}\left({W*m}_{Graph}^{T}\right)\oplus{Emb}_{spatial}\left({W*m}_{Graph}^{T}\right)\oplus{Emb}_{time}\left({W*m}_{Graph}^{T}\right)$, where $W\in R^{{hidden}*d}$ is the weight of the linear layer.
Oblique Photography: The oblique photography modality consists of images from multiple views, i.e., $m_{Oblique}=\{V_{1}\in R^{C*H*W},\ V_{3}\in R^{C*H*W},V_{3}\in R^{C*H*W},\ldots\ ,V_{n}\in R^{C*H*W}\}$, where $V_{i}$ represents the image of each view and n is the number of views. In the encoder for the oblique photography modality, we use the standard ViT model to extract features from multiple view images and concatenate them. The process can be formalized as $f\left(m_{Oblique}\right)={Enc}_{12}\left(V_{1}\right)\oplus{Enc}_{12}\left(V_{2}\right)\oplus{Enc}_{12}\left(V_{3}\right)\oplus\ldots\oplus{Enc}_{12}\left(V_{n}\right)$.
@@
Video: The video modality can be considered as the stacking of images in the temporal dimension, i.e., $m_{video}\in R^{T*C*H*W}$. The significant characteristic of the video modality is its high redundancy in both temporal and spatial dimensions. To address this issue, the encoder for the video modality in AllSpark is based on TubeViT[40], which performs sparse sampling on highly redundant data in both time and space to reduce redundancy. The features obtained from sparse sampling are subsequently fed into a 6-layer transformer encoder for further feature extraction. The process can be formalized as $f\left(m_{video}\right)={Enc}_{6}\left({Sparse}(m_{video})\right)$.
PointClouds: PointClouds are typically composed of three-dimensional coordinates and feature values $m_{PointCloud}\in R^{K*\left(d+3\right)}$, where K represents the number of three-dimensional points and d represents the dimensionality of the point cloud features, reflecting information such as the spatial position, shape, colour, and texture of objects. The encoder for the point cloud modality inherits the design from PointBERT[41]: first, point cloud data are grouped and encoded to unify the number of points simultaneously inputted. This step can be represented as $PointGroup=Grouper\left(m_{PointCloud}\right)\in R^{G*N*3}$, where $G$ represents the number of groups and $N$ represents the number of points in each group. Next, the grouped results are input into a one-dimensional convolutional layer to extract feature vectors for each group:$f_{Group}=Conv1d\left(PointGroup\right)\in R^{G*d}$. Finally, the feature vectors for each group are input into a standard 12-layer transformer encoder to extract their global features. The entire process can be formalized as $f\left(m_{PointCloud}\right)=Enc_{12}\left(Conv1d(Grouper(m_{PointCloud}))\right)$.
@@ III-D Modal2Language Bridge
Although the modal encoders have transformed data from various heterogeneous modalities into a unified token sequence, there are still differences in dimensions between different modal tokens, making it difficult to perceive by a multimodal large language model. The modal bridge, based on the Lynx model, aims to perform dimensional projection from tokens of various modalities to tokens of the language modality. In its implementation, the modal bridge consists of stacked layers, including cross-attention layers and feedforward neural network layers.
@@
The feedforward neural network inherits the classic design from the original transformer and consists of two linear layers with an inserted activation layer. The entire process can be formalized as follows:
@@ III-E Task-Guided Text Prompts and Task Heads
To extend the visual-language multimodal model to thirteen spatio-temporal modalities without intervention from modality expert knowledge, we designed specific text prompts and task heads for each modality and task. Text prompts are used to guide the multimodal language model in correctly interpreting each modality’s data, while task heads are employed to match the model’s parsing results with specific downstream tasks.
We manually designed one to four specific text prompts for each modality. During the training process, to enhance model performance, we employed a strategy of diversifying prompts, randomly selecting one prompt for each forward pass. However, during testing, for the sake of result stability and reproducibility, the prompt was fixed to be the first prompt among all prompts. Table XVI provides a list of all the text prompts.
Additionally, we noticed that there is still a gap between the multimodal language model’s parsing results and the specific downstream tasks. To enhance the model’s generalization capability on downstream tasks, we designed specific task heads for each task. Moreover, to ensure the transferability of modalities across different tasks, the design principle for task heads is to be as simple and lightweight as possible. A detailed summary of each modality’s task heads can be found in Table 4.
For classification tasks or downstream tasks that can be formalized as classification tasks, we uniformly use a simple single-layer linear layer as the task head. For instance, we implemented standard classification tasks on the RGB, MSI, SAR, video, and PointCloud modalities. Although the task involving the HSI modality is segmentation, it can be formalized as a per-pixel classification task. Similarly, the reidentification task in the infrared modality can also be formalized as a classification task. Therefore, a single-layer linear layer is used as the task head for the mentioned modalities.
For regression tasks on the table, trajectory, and graph modalities, we also use a linear layer to perform regression predictions. The only difference from classification tasks is the addition of an unscaled operation without learnable parameters. Since the Lynx model itself is a language model, the code and text modalities directly use its native text decoder. The task head for the oblique photography modality is based on the 3D reconstruction architecture from the Ada-MVS[42] model.
@@ IV-A Setup
TABLE II: Summary of hyperparameters
Dimension
Modal
Max Lr
Max Epochs
Warm-up Epochs
1D
Text
$9.0\times 10^{6}$
5
1
Code
$1.0\times 10^{5}$
4
1
2D
RGB
$5.0\times 10^{5}$
50
5
MSI
$2.0\times 10^{5}$
50
5
HSI
$1.0\times 10^{4}$
30
3
Infrared
$5.0\times 10^{5}$
50
5
SAR
$9.0\times 10^{6}$
30
3
Oblique Photography
$5.0\times 10^{5}$
30
3
Table
$2.0\times 10^{5}$
30
3
Trajectory
$1.0\times 10^{5}$
30
5
Graph
$8.0\times 10^{5}$
10
2
3D
PointCloud
$3.0\times 10^{5}$
100
10
Video
$1.0\times 10^{5}$
3
1
Our experiments evaluated the adaptability of AllSpark to various spatio-temporal modalities. To provide a realistic evaluation of its multimodal adaptability and facilitate comparisons with existing work in the field, we followed the principles of simplicity, reproducibility, and choosing datasets and tasks that have been widely studied in the domain. The specific details on dataset selection can be found in Section IV-B.
Additionally, for reproducibility, we employed similar experimental settings across all modalities. We used the AdamW optimizer with a learning rate schedule based on cosine annealing. The hyperparameters were adjusted slightly in terms of training epochs and learning rates for different experiments. Table II summarizes the hyperparameter settings for the experiments on each modality.

### v3

@@ III-A Language as Reference Framework
Inspired by our observation, we introduce the fundamental principle of Language as Reference Framework (LaRF) to balance cohesion and autonomy among multiple modalities.
Figure 2: Guided by the LaRF principle, multimodal data are transformed into a token-context structure akin to language, based on their respective prior assumptions. This approach preserves the autonomy of each modality while achieving cohesion between them, enabling the interpretation of multimodal data within a unified language representation space.
In terms of the cohesion of multimodalities, the high heterogeneity between multiple spatio-temporal modalities is a major challenge, while the LaRF principle defines the alignment anchor between multimodalities as language explicitly. As shown in Figure 2, we observe that language is encoded by tokens and their contexts, and this structure can be extended to most spatio-temporal modalities. Therefore, we can align highly heterogeneous spatio-temporal modalities to language modalities in structure and semantics, enabling multimodal interpretation in the unified representation space of language. Additionally, the pivotal role of natural language prompts is a key factor in the LaRF principle’s ability to achieve cohesion across modalities.
@@
    Interpretability: Deep learning methods have long been characterized as ”black boxes.” However, a multimodal intelligent system constructed based on the LaRF can directly leverage language as a tool. This facilitates the output of interpretable reasoning chains that humans can understand, thereby achieving true explainable artificial intelligence.
@@
Figure 3: AllSpark Architecture. Multimodal data are extracted by their respective modal encoders into token sequences. Following dimension alignment with modality-specific text prompt tokens via a modal bridge, both the text prompt tokens and modality tokens are passed into a large language multimodal model for interpretation. The interpretation results are then aligned with downstream tasks through task-specific heads.
@@ III-B Overview
TABLE II: List of symbols
@@
Symbol Meaning Symbol Meaning
$M$ AllSpark $m_{i}$ modal i
$f_{i}$ modal encoder for $m_{i}$ $q$ query vectors of the bridge
$\Phi$ modal bridge $s_{i}$ tokens of $m_{i}$
$T$ text tokenizer $p_{i}$ text prompt for $m_{i}$
$F$ Multimodal LLM $\theta$ model parameters
$H_{task}$ Task Head $y$ label
$L$ loss function $t_{i}$ tokens of $m_{i}$
${Enc}_{n}$ n-layer transformer encoders $W$ weights of the linear layer
$Emb$ embedding layer $w_{i}$ word of the code or text
${Enc}_{ResNet}$ ResNet
$Grouper$ point grouper of the PointBERT[48]
$Conv1d$ 1D convolution layer
$FFN$ feedforward network
$\sigma$ softmax
Guided by the principles mentioned above, AllSpark consists of five modules: the modal encoder, modal bridge, text tokenizer, multimodal large language model, and task head. The overall architecture is depicted in Figure 3.
@@
Here, $\{m_{RGB},m_{MSI},m_{HSI},\ldots,m_{i},\ldots\}$ represents inputs from various modalities, where $m_{RGB}$ represents the RGB modality, $m_{MSI}$ represents the multispectral modality, and so on. $f_{i}$ is defined as the modal encoder for the $m_{i}$ modality, $\Phi$ represents the modal bridge, and $q\in R^{N*D}$represents $N$ learnable vectors of dimension $D$ in the modality bridge, where $D$ is set to 4096, representing the dimensionality of the multimodal large language model. The input data $m_{i}$ of each modality are mapped to a token sequence $s_{i}\in R^{N*D}$ of the same dimension as the language model. Table II summarizes the main mathematical symbols and their meanings in the model.
To achieve cohesion among modalities, we employ a unified multimodal large language model to parse data from various modalities. The text tokenizer and multimodal large language model in AllSpark are based on the visual-language model Lynx[49]. To extend Lynx to ten spatio-temporal modalities, we designed specific text prompts for each modality to guide the model in correctly parsing information from each modality. Additionally, Lynx incorporates several lightweight multimodal adapter layers internally to accommodate multimodal inputs. We continue this design and do not freeze the parameters of the adapter layers during training to enhance the model’s adaptability to other spatio-temporal modalities. Finally, we acknowledge that a gap exists between the parsing results of the model and those of the downstream task. Therefore, we design specific task heads for each task to enhance the model’s generalization capability.
@@ III-C Independent Encoder for Each Modality
Code & Language: Code is essentially a specialized form of language. But due to its distinct properties such as having a strict syntax, being unambiguous, and being capable of interacting with machines, we separate it as a distinct modality. Therefore, to avoid ambiguity, we will use ”text” and ”language” interchangeably to distinguish between the natural language and code modalities. Exploring intelligent methods for the code modality is crucial for reliable AI reasoning and achieving interaction between intelligent models and the real world. Since the Lynx model is a language model, we do not design an additional modal encoder for the code and text modalities. Instead, we directly utilize Lynx’s text tokenizer, i.e., $f\left(m_{Text/Code}\right)=T\left(m_{Text/Code}\right)$, where $m_{Text/Code}=\{w_{1},w_{2},w_{3},\ldots\}$ represents the sequence of words in the text or code.
Table: A table can be viewed as a sequence of rows, with each row containing several fields, or columns, i.e., $m_{Table}\in R^{{row}*{col}}$. The modal encoder for the table modality inherits the design from TabFormer[50]: first, based on the different degrees of discreteness for each column attribute, we use independent embedding layers to encode discrete and continuous values separately. Subsequently, we employ a single-layer transformer encoder to further extract features. The entire process can be formalized as $f\left(m_{Table}\right)={Enc}_{1}\left(Emb(m_{Table})\right)$.
@@
RGB: RGB imagery represents the visible light spectrum and reflects the electromagnetic characteristics of objects that emit or reflect visible light waves. It is the most common modality in the field of computer vision. RGB imagery is a standard three-band image, i.e., $m_{RGB}\in R^{H*W*3}$. For the modal encoder of this modality, we adopted the visual encoder from the Lynx model: EVA[51]. The EVA is a large visual model composed of 40 stacked transformer blocks with a width of 1408. During the experiments, AllSpark loaded the official weights of the EVA model and froze them during training.
MSI: Multispectral imagery is a modality extensively studied in the remote sensing field. It incorporate multiple nonvisible light bands, such as near-infrared, shortwave infrared, coastal atmospheric aerosol, and cirrus bands. Therefore, the number of channels in MSIs is usually greater than the three bands in RGB imagery, i.e., $m_{MSI}\in R^{H*W*C}$, where $C>3$. We extended the PatchEmbed of the standard ViT[31], modifying its channel count to match the number of bands in the input multispectral imagery. This adaptation allows it to serve as the feature encoder for the multispectral modality.
@@
Graph: A graph is constructed from a series of nodes and edges, where the attributes of the nodes and the adjacency attributes of the nodes reflect the majority of the features of the graph, i.e., $m_{Graph}\in R^{K*d}$, where K is the number of nodes and d is the node feature dimension. In AllSpark, the modality encoder of a graph is based on the STAEformer[53, 54], whose main design idea is to first use a linear layer to extend its feature dimension and then use several embedding layers to separately encode features such as the node’s characteristics, spatial characteristics, and temporal characteristics. The entire process can be formalized as $f\left(m_{Graph}\right)={Emb}_{node}\left({W*m}_{Graph}^{T}\right)\oplus{Emb}_{spatial}\left({W*m}_{Graph}^{T}\right)\oplus{Emb}_{time}\left({W*m}_{Graph}^{T}\right)$, where $W\in R^{{hidden}*d}$ is the weight of the linear layer.
@@
Point Cloud: A point cloud is typically composed of three-dimensional coordinates and feature values $m_{PointCloud}\in R^{K*\left(d+3\right)}$, where K represents the number of three-dimensional points and d represents the dimensionality of the point cloud features, reflecting information such as the spatial position, shape, colour, and texture of objects. The encoder for the point cloud modality inherits the design from PointBERT[48]: first, point cloud data are grouped and encoded to unify the number of points simultaneously inputted. This step can be represented as $PointGroup=Grouper\left(m_{PointCloud}\right)\in R^{G*N*3}$, where $G$ represents the number of groups and $N$ represents the number of points in each group. Next, the grouped results are input into a one-dimensional convolutional layer to extract feature vectors for each group:$f_{Group}=Conv1d\left(PointGroup\right)\in R^{G*d}$. Finally, the feature vectors for each group are input into a standard 12-layer transformer encoder to extract their global features. The entire process can be formalized as $f\left(m_{PointCloud}\right)=Enc_{12}\left(Conv1d(Grouper(m_{PointCloud}))\right)$.
@@ III-D Modal2Language Bridge
Although the modal encoders have transformed data from various heterogeneous modalities into a unified token sequence, there are still differences in dimensions between different modal tokens, making it difficult to perceive by a multimodal large language model. The modal bridge, based on the Perceiver[21], aims to perform dimensional projection from tokens of various modalities to tokens of the language modality. In its implementation, the modal bridge consists of stacked cross-attention layers and feedforward neural network layers.
@@
The feedforward neural network inherits the classic design from the original transformer and consists of two linear layers with an inserted activation layer.
The entire process can be formalized as follows:
@@ III-E Task-Guided Text Prompts and Task Heads
To extend the visual-language multimodal model to ten spatio-temporal modalities without intervention from modality expert knowledge, we designed specific text prompts and task heads for each modality and task. Text prompts are used to guide the multimodal language model in correctly interpreting each modality’s data, while task heads are employed to match the model’s parsing results with specific downstream tasks.
We manually designed one to four specific text prompts for each modality. During the training process, to enhance model performance, we employed a strategy of diversifying prompts, randomly selecting one prompt for each forward pass. However, during testing, for the sake of result stability and reproducibility, the prompt was fixed to be the first prompt among all prompts. Table XVII provides a list of all the text prompts.
To ensure the transferability of modalities across different tasks, the design principle for task heads is to be as simple and lightweight as possible.
For classification tasks or downstream tasks that can be formalized as classification tasks, we uniformly use a simple single-layer linear layer as the task head. For instance, we implemented standard classification tasks on the RGB, MSI, SAR, and point cloud modalities. Although the task involving the HSI modality is segmentation, it can be formalized as a per-pixel classification task. Therefore, a single-layer linear layer is used as the task head for the mentioned modalities.
For regression tasks on the table, trajectory, and graph modalities, we also use a linear layer to perform regression predictions. The only difference from classification tasks is the addition of an unscaled operation without learnable parameters. Since the Lynx itself is a language model, the code and text modalities directly use its native text decoder.
TABLE III: Summary of hyperparameters
Dimension Modal Max Lr Max Epochs Warm-up Epochs
1D Language $9.0\times 10^{6}$ 5 1
Code $1.0\times 10^{5}$ 4 1
Table $2.0\times 10^{5}$ 30 3
2D RGB $5.0\times 10^{5}$ 50 5
MSI $2.0\times 10^{5}$ 50 5
HSI $1.0\times 10^{4}$ 30 3
SAR $9.0\times 10^{6}$ 30 3
Trajectory $1.0\times 10^{5}$ 30 5
Graph $8.0\times 10^{5}$ 10 2
3D Point Cloud $3.0\times 10^{5}$ 100 10
@@ IV-A Setup
Our experiments aim to demonstrate: (1) the distinct advantages of AllSpark over traditional models, and (2) AllSpark’s ability to understand 10 spatio-temporal modalities simultaneously.
For the former, we believe that since language as reference framework, AllSpark’s advantage over traditional models lies in the richer semantic meaning of its features. In large language models[55, 26, 27, 56], a common way to measure this property is through few-shot learning. Therefore, we evaluated AllSpark’s performance on few-shot classification tasks in the RGB and point cloud modalities without any extra training steps. The experimental results can be found in Section IV-C.
For the latter, we select a task for each modality to conduct the evaluation. Following the principles of simplicity and reproducibility, we choose the widely studied datasets in each modality’s respective field and employ similar experimental settings across all modalities. Specifically, we use the AdamW optimizer with a learning rate schedule based on cosine annealing. The hyperparameters are adjusted slightly in terms of training epochs and learning rates for different experiments. The specific details on dataset selection can be found in Section IV-B. Table III summarizes the hyperparameter settings for the experiments on each modality. The experimental results can be found in Section IV-D.